# Light Future: Multimodal Action Frame Prediction via InstructPix2Pix

### v1

@@ 4.3 Model Fine-tuning Strategy
Progressive Training Strategy: We employ a two-stage training method: first fine-tuning the model at a lower resolution (64$\times$64), then gradually increasing to the target resolution (128$\times$128). This strategy helps stabilize the training process and improve final performance.
@@ 5.1 Dataset
We used the RoboTwin simulator to generate our training and testing dataset. For each sample in the dataset, it contains approximately 400 frames, covering an initial frame, a text instruction, and a ground truth frame 100 steps later. we focused on three specific robotic tasks:
@@
In the first stage of experiment, for each task, we generated 100 observations, resulting in a total of 300 samples. In the second stage of experiment, we constructed the training set by pairing every 10th frame from each sample (e.g., frame 0 - frame 100, frame 10 - frame 110, …), ultimately creating a total of 10,491 image pair samples.
@@ 5.4 Results
Our model achieves high-quality future frame predictions. In stage 1, SSIM improved from 0.9391 to 0.9794, and PSNR increased from 54.30 to 59.19 as training epochs increased from 10 to 50. In stage 2 of the experiment, with more diverse samples and broader range of robotic motions, only 2–10 epochs were required to reach comparable and high performance (SSIM: 0.9823, PSNR: 59.41).
@@
10 epochs 0.9391 54.30
50 epochs 0.9794 59.19
Table 2: Quantitative evaluation of our model’s performance on test set in the first stage of experiment.
@@
2 epochs 0.9766 58.04
10 epochs 0.9823 59.41
Table 3: Quantitative evaluation in the second stage of experiment.
@@ 5.5 Qualitative Analysis
Figure 7 illustrates the predicted consecutive multiple frames and consecutive ground truth images (we randomly selected the 47th frame as the input, and we could see the predicted results after 100 frames (10 seconds)). The results demonstrate that our approach generates highly accurate action trajectory predictions, with the forecasted motion closely aligning with the ground truth (image fidelity is not our primary focus). Experimental validation confirms that robust trajectory prediction with our framework can be achieved with just 10,000 training pairs and only 10 epochs of training.
@@ 5.6 Inference time for predicting future frames
Table 5 demonstrates the inference time requirements of mainstream video prediction models. Our design exhibits significant advantages in lightweight architecture and inference speed, particularly when compared to the computationally intensive nature of multimodal models.
Without applying any specific optimizations for inference acceleration, our design currently just requires 38 seconds to generate 15 future frames, which is much faster than many other ”heavy” models like Flamingo-3B in application realm. We identify significant opportunities for speedup through subsequent optimizations such as quantization and other inference acceleration techniques. Our framework’s efficiency advantage proves especially impactful in action prediction scenarios.
Table 5: Comparison of Video Prediction Models (15-frame generation)
Model Parameters Inference Time
VDM 2B 6–8 min
LVD 3B 10–12 min
Flamingo-3B 3B 15+ min
MAGVITv2 1.2B 4–6 min
MVCD 1.5B 6 min
InstructPix2Pix (FT) 1.5B 38 sec
@@ 5.9 Ablation Study
To understand the contribution of different components in our approach, we conducted a simple ablation study by varying the number of training epochs while keeping other hyperparameters constant. When increasing the number of epochs from 2 to 10 led to significant improvements in both SSIM (+0.0057) and PSNR (+1.37dB). This highlights the importance of sufficient training iterations for the model to capture the nuances of robotic movements and accurately predict future frames. In our second stage of the experiment, we trained for 10 epochs, and we believe that as the number of epochs continues to increase, we anticipate seeing continued performance gains.
@@ 6 Conclusion
In this paper, we have successfully implemented and evaluated a deep learning-based approach for robotic action visual prediction. By fine-tuning the InstructPix2Pix model on RoboTwin simulation data, we developed a system capable of predicting what a robot will see multiple frames after executing a specific action.
This design transforms the computationally expensive video frame prediction task into a more lightweight and controllable multimodal vision-language controlled prediction task. This approach not only improves operational efficiency but also achieves fine-tuning and future frame inference with significantly reduced GPU resource requirements. Furthermore, our experimental results demonstrate the effectiveness of this approach, achieving high-quality predictions with SSIM values of 0.9823 and PSNR of 59.41dB in robot action prediction task.

### v2

@@ 4.3 Model Fine-tuning Strategy
Progressive Training Strategy: We employ a two-stage training method: first fine-tuning the model at a lower resolution ($\times64$), then gradually increasing to the target resolution ($\times128$). This strategy helps stabilize the training process and improve final performance.
@@ 5.1 Dataset
We used the RoboTwin simulator to generate our training and testing dataset. For each sample in the dataset, it contains approximately 400 frames, covering an initial frame, a text instruction, and a ground truth frame 100 steps later. we focus on three specific robotic tasks:
@@
In the first stage of the experiment, for each task, we generated 100 observations, resulting in a total of 300 samples. In the second stage of the experiment, we constructed the training set by pairing every 10th frame from each sample (e.g., frame 0 - frame 100, frame 10 - frame 110, …), ultimately creating a total of 10,491 image pair samples.
@@ 5.4 Results
Our model achieves high-quality future frame predictions. In stage 1, SSIM improved from 0.88 to 0.90, and PSNR increased from 36.30 to 38.19 as training epochs increased from 10 to 50. In stage 2 of the experiment, with more diverse samples and broader range of robotic motions, only 2–10 epochs were required to reach comparable and high performance (SSIM: 0.93, PSNR: 39.71).
Table 2: Quantitative evaluation of our model’s performance on test set in the first stage of the experiment.
@@
10 epochs 0.8871 36.30
50 epochs 0.9094 38.19
Table 3: Quantitative evaluation in the second stage of the experiment.
@@
2 epochs 0.9066 38.04
10 epochs 0.9323 39.71
@@ 5.5 Qualitative Analysis
Figure 7 illustrates the predicted consecutive multiple frames and consecutive ground truth images (we randomly selected the 47th frame as the input, and we could see the predicted results after 100 frames (10 seconds)). The results demonstrate that our approach generates highly accurate action trajectory predictions, with the forecasted motion closely aligning with the ground truth (In scenarios such as robotic movement trajectory, precise prediction of motion trajectories is our primary concern, while image fidelity is not the foremost priority). Experimental validation confirms that robust trajectory prediction with our framework can be achieved with just 10,000 training pairs and only 2-10 epochs of training.
@@ 5.6 Inference Time for Predicting Future Frames
In our experiment on an NVIDIA A100 (40GB), our fine-tuned InstructPix2Pix generates 15 future frames in about 38 seconds. For reference, large multimodal baselines typically require minutes under similar conditions: Flamingo-3B takes over 12 minutes, while LVD needs around 8–10 minutes. This highlights the efficiency advantage of our lightweight adaptation for robotic prediction tasks. Even without specific inference optimizations, our framework shows clear efficiency advantages for robotic prediction tasks, with further opportunities for speedup via quantization and other acceleration techniques.
@@ 5.9 Ablation Study
To understand the contribution of different components in our approach, we conducted a simple ablation study by varying the number of training epochs while keeping other hyperparameters constant. When increasing the number of epochs from 2 to 10 led to significant improvements in both SSIM and PSNR. This highlights the importance of sufficient training iterations for the model to capture the nuances of robotic movements and accurately predict future frames. In our second stage of the experiment, we trained for 10 epochs, and we believe that as the number of epochs continues to increase, we anticipate seeing continued performance gains.
@@ 6 Conclusion
In this paper, we present a deep learning framework for robotic action visual prediction. By fine-tuning the InstructPix2Pix model on RoboTwin simulation data and create a new framework, our method predicts the robot’s future observation frames from a current action frame and a textual instruction.
This design transforms the computationally expensive video frame prediction task into a more lightweight and controllable multimodal vision-language controlled prediction task. This approach not only improves operational efficiency but also achieves fine-tuning and future frame inference with significantly reduced GPU resource requirements. Furthermore, our experimental results demonstrate the effectiveness of this approach, achieving high-quality predictions with SSIM values of 0.93 and PSNR of 39.71dB in robot action prediction task.